Could someone help me out here; find the coordinates of the point C, halfway between the points A(5,1) and B(-2,7). I dont really understand how this works =(
Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(69443)

find the coordinates of the point C,
halfway between the points A(5,1) and B(-2,7).
------
Find the midpoints of x and y separately by getting the averages.
for x: (5-2)/2 = 3/2
for y: (1+7)/2 = 4
C is (3/2,4)

Answer by scott8148(6628)

the midpoint of two sets of coordinates (points) is halfway between the two x-coordinates and halfway between the two y-coordinates

taking the average of the x's and the average of the y's will give you the midpoint

midpoint (M) is [(5-2)/2 , (1+7)/2) ___ M(1.5, 4)